4a) Determine the five constitutional isomers for N2O.  Three have the N-N-O bond skeleton and two would have the N-O-N skeleton. Determine the formal charges on all atoms of each of the isomers (write # on molecules above). (0.5 pt)

b) Write the net formal charge of each isomer.  Label molecules above (1-5) (0.25 pt)

1. ______        2. ______        3. ______        4. ______        5. ______

c) Which isomer is most stable?  Why? (0.25 pt)

6. Of the two type of bonds (pi and sigma), which is capable of rotation around that bond and which is not? Why?
Answer #1

Answer:- the five constitutional isomers of N2O is given below:-

b):- Formula for calculating formal charge ( F) = [ no. of valance electrons on atom in free state] - [ no.of non- bonding electrons] - 1/2 [ no. Of bonding electrons]

Or

F = V - (NB) - 1/2 B

​​​​​---------------------

Formal charge on structure 1 atoms( N=N=O):-

Formal charge on N# atom = 5- 4-(1/2 × 4) = 5-4-2 = -1

Formal charge on N## atom = 5- 0 - (1/2 × 8) = 5-4 = +1

Formal charge on O atom = 6- 4 - (1/2 × 4) = 6 -4 -2 = 0

------------------

Formal charge on structure 2 atoms,

Formal charge on N# atom = 5-2-(1/2 × 6) = 5-2-3 = 0

Formal charge on N## atom = 5-0 - 4 = +1

Formal charge on O atom = 6- 6 - 1 = -1

----------------

Formal charge on structure 3 atoms,

Formal charge on N# atom = 5-6-1 = -2

Formal charge on N## atom = 5-0-4 = +1

Formal charge on O atom = 6-2-3 =+1

--------------

Formal charge on structure 4,

Formal charge on N# atom= 5-6-1 = -2

Formal charge on O atom = 6-0-4 = +2

Formal charge on N## atom = 5- 2-3 =0

-----------------

Formal charge on structure 5,

Formal charge on N# atom= 5-4-2 = -1

Formal charge on O atom = 6-0-4 = +2

Formal charge on N## atom = 5- 4-2 = -1

----------------------

c) :-

According to formal charge concept, a stable structure of a molecule is one with less no. of formal charges on its atoms. Accordingly, both structures 1 & 2 are stable, as they contain least formal charges on their atoms. But structure 1, is most stable because of presence of negative charge on oxygen atom as compared to negative charge on N atom on structure 2. We all know , that more electronegative atom ( O atom) stabilizes -ve charge than less electronegative atom (N atom).

---------------

6. Sigma bond is capable of rotation , while pi bond restrict rotation. This is because of overlapping of pi- orbitals above and below the plane of the atoms.